Fermi National Laboratory


Accelerator Update

Wednesday November 26
The day shift began with Operations monitoring store 3048, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:19 AM, Operations terminated store 3048. Unfortunately, the Tevatron (TeV) suffered a sector A1 quench on termination.

A high voltage expert reported that technicians had the Meson feeder 30 powered by 9:30 AM.

At 9:48 AM, a Preacc expert removed the I- Source for repairs. This caused the Linac buncher to stop working. Beam resumed about an hour later.

Operations put the TeV into shot setup at noon. At 2:49 PM, Operations established store 3050 with an initial luminosity of 1.57E31. Pbar resumed stacking.

Two Booster RF stations (BRF5 & 6) tripped off at 5 PM due to a vacuum burst. They both reset okay.

Thursday November 27
The midnight shift began with Operations monitoring store 3050, with Pbar stacking, and with MiniBooNE taking beam.

At 6:50 AM, Operations received a LCW leak alarm for the Main Injector (MI). Operators accessed MI and discovered the leak. Here is what they saw:


MI-60


The Results of a Blown Hose

The Hose

There were actually three leaks eventually found that caused the MI tunnel to turn into the MI canal: a blown hose and bad valve in MI, and a blown hose in MiniBooNE. Operators valved off the leaks and then, with help from the mechanical techs, began flood control and clean up. Water group techs came in and began repairs. There were many devices that were soaked and needed drying off before beam could again go through MI, and this was a slow process since not all of these leaks were found at the same time.

By 4 PM, everyone was out of MI and Operations had begun a hipot and final checkout. The hipot didn't look good.

At 8:30 PM Operators found the second MI leak on a Lambertson magnet. Water techs again came in and made repairs.

Operators found the third leak at 10:38 PM, at sector MI12B. Operators valved off the hose. (An expert came in on Friday morning to make repairs.)

Friday November 28
The midnight shift began with MI off due to a LCW repairs and cleanup, with Operations monitoring store 3050,

Operations ran Switchyard and Meson power supplies from around 4 AM to 7:30 AM.

At 6 AM, the TeV aborted and suffered a sector A1 quench. A TeV RF station (TRF2) had tripped, which caused the abort.

Operations turned on MI around 7 AM and resumed Pbar stacking at 7:18 AM.

At 10:22 AM, Operations put the TeV into shot setup and at 12:46 PM, established store 3052. Its initial luminosity was 2.21E31.

Operation resumed sending beam to MiniBooNE around 2 PM.

At 6 PM, Operations sent beam into the Recycler for studies.

The MiniBooNE Horn tripped for the first of several times due to a water problem at 9 PM. Operators contacted a Horn expert.

Saturday November 29
The midnight shift began with Operations monitoring store 3052, with the MiniBooNE horn off due to water problems, and with Pbar stacking.

Operations reestablished beam to MiniBooNE at 2 AM.

At 11 AM, Operations terminated store 3052. The TeV quenched at sector A1 on termination. Techs changed out a lead flow restrictor in sector E2.

MiniBooNE was halted due to airflow interlock problems. An interlock tech came in and made adjustments. Beam to MiniBooNE resumed around 3:30 PM.

Operations put the TeV into shot setup and established store 3053 at 5:10 PM. Its initial luminosity was 1.8E31. Pbar resumed stacking.

At 5:40 PM, experts conducted Booster aperture scans. This held off MiniBooNE and caused the stacking rate also suffered due to these scans. Everything was back to normal by 8:30 PM.

Sunday November 30
The midnight shift began with Operations monitoring store 3053, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler experts completed their studies at 1:30 PM.

At 5:19 AM, a Linac expert began a 400 MeV area aperture scan. This again shut off beam to MiniBooNE and stacking until 12:23 PM, when the expert completed his scans.

Operations terminated store 3053 at 1:10 PM. The TeV suffered a sector A1 quench on termination.

At 2:30 PM, the Pbar emittance began growing large. Operators cooled the stack. This problem continued until 6:30 PM.

A problem with a Linac RF station (LRF1) required operators to call in a tech at 4:41 PM. He adjusted the filament current and had the station regulating better by 5:30 PM.

A Linac expert began a half-hour aperture scan at 8:30 PM.

At 8:50 PM, the H- Source came into alarm due to a gas pressure problem. Operators reset a vacuum pump to resolve the problem.

Operations put the TeV into shot setup at 9:37 PM.

Monday December 1
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established store 3057 at 12:46 AM, with an initial luminosity of 2.2E31.

Pbar experts accessed their tunnels at 1 AM.

Linac buncher problems began at 1:47 PM. Operator kept making adjustments until 3 AM, when they called in an expert. At 4:30 AM, the expert reported that the temperature in the 750 KeV line was colder than normal and that was affecting the buncher. He found a simple fix to the problem by covering the buncher. Beam resumed at 5:08 AM.

Today's Plans
The plan for today is to have TeV experts investigate the quench on termination problem and then establish a store later this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/1/2003   email Fermilab